Project

General

Profile

20170523 » History » Revision 5

Revision 4 (Boyd, Steven, 22 May 2017 22:07) → Revision 5/11 (Rajaram, Durga, 23 May 2017 13:10)

h1. Ops/Comm Scheduling meeting : 23/5/2017 

 Location : CR21 at 14:00 pm 

 Phone access :    *New* phonebridge number 
 UK: 08082380274 
 US: 631.621.5253 
 Access code: 1779339517 

 h2. Agenda 

 * Actions 
 ** -Macwaters - Re-make dodgy compressor2 remote cable connector - stands- 
 ** Macwaters - Remove old overheated CSW71 to R9 _waiting for John Govans and the van_ - in progress 
 ** -Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress- 
 ** Macwaters - Report on status of FC psu repair (stands) 
 *** The spare is installed and has been run into a short  
 ** Hall Manager - water system change-over procedure (stands) - _collar people on Friday_ 
 ** EOverton - Status of @miceraid5@ disk cleanup (stands) _now has a script but needs to run it_ 
 ** Progress on Hall list    [[Hall_Task_List_Spring17|List of Hall activities in the shutdown]] (Colin) _DL were down on masse and made significant progress. Need to confirm if any remaining items_ 
 ** CWhyte to request survey of TOF1 - _requested, next week as Tony off this week_ 
 ** SBoyd to ensure we get updates on detector/data quality and on Target performance at Ops meetings 
 ** AKurup to check which settings are being used by the active Alarm Handlers 
 ** SBoyd to conscript extra shift cover for the end of next week : Done. Need BLOCs for the 31/1 and 2 of June (i.e. end of cycle) 

 * DC update 
 * MOM update 
 * Performance plots from TOF calibration run (Durga?) attachment:Onrec-9297.pdf 
 * Update on the tracker saga 
 * Hydrogen system update 
 * Run plan 
 ** Tracker may be ready to operate on Friday 
 *** Is there still Helium in the tracker? 
 ** LH2 running will not happen this cycle 
 ** Extraction of LH2 absorber, re-insertion of the FC and establishment of the magnetic channel: Thursday (25May17)—Saturday (27May17); 
 ** Magnet commissioning: Sunday (28May17)—Monday (29May17); and 
 ** Empty-channel data taking (Monday 29May17)—Friday (02Jun17). 
 ** Last training date on Friday. I've asked shifters to meet the MOM on Friday to assess training plans. The MOM should set a time for this meeting. 
 * AOB 

 * Last meeting : [[20170516|Tuesday May 9 2017]] 
 * Next meeting : [[20170530|Tuesday May 30 2017]]